Circular No. 7476

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000dc IN ESO 527-G019 C. Yu and W. D. Li, University of California at Berkeley, on behalf of

the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Aug. 9.3 UT (mag ∼ 16.9) and Aug. 10.3 (mag ∼ 16.5). The new object is located at α = 20^{h}20^{m}45^{s}.47, δ = −24^{o}07′57′′.5 (equinox 2000.0), which is 3′′.5 west and 16′′.0 south of the nucleus of ESO 527-G019. A KAIT image of the same field on Aug. 2.3 showed nothing at the position of SN 2000dc (limiting mag ∼ 19.0).

COMET C/1999 S4 (LINEAR) H. Weaver, Johns Hopkins University, and R. West, European South-

ern Observatory, on behalf of a large group of collaborators, report the following results: “Hubble Space Telescope (HST) images taken during Aug. 5.167–5.396 UT and Very Large Telescope (VLT) images (image quality ∼ 0′′.6) taken during Aug. 6.978–6.999 revealed about a dozen active fragments, most of them located within ∼ 20′′ of the western tip of the dust tail (cf. IAUC 7474). The correspondence between fragments in the HST and VLT images is generally very good, but the brightest fragment in the HST image is not seen in the VLT image, indicating rapid variability in the activity levels. The dynamic nature of the fragments was further highlighted by a dramatic change in the appearance of the fragments in VLT images taken during Aug. 9.976–9.996, when they were barely detectable. Although the latter images were taken under mediocre observational conditions (image quality ∼ 1′′–1′′.3, thin cirrus, and nearly full moon), that alone seems unlikely to account for the observed changes. A very preliminary estimate for the R magnitude, within a circular aperture of radius 0′′.23, of one well-isolated fragment is ∼ 24. A completely inactive fragment with a diameter of 100 meters observed under these conditions (r = 0.79 AU, Δ = 0.69 AU, β = 86^{o}) would have R ∼ 25.9 (assuming a 4-percent albedo and 0.04 mag/deg phase law). We urge ground-based observers to continue monitoring the comet and to report any unusual changes near the ‘tip’, both in brightness and morphology.”

2000 August 10 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT Elizabeth O. Waagen